# Freeze-in Gravitational Waves and Dark Matter in Warm Inflation

### v3

@@ 4 Gravitational waves spectrum
withk^=k/T\ $\hat${k}=k/T. Here $k$ represents the momentum of graviton. The $\hat{\eta}$ function takes into account all the Standard Model contributions and has been calculated in refs. [22, 23] using the leading-log approximation. The calculation including hard thermal loop resummation has been done in ref. [24]. Here we only consider the results of leading-log approximation [22],
@@
$\rho_{\mathrm{GW}}(N_{e})=\frac{2}{\pi^{2}M_{\rm pl}^{2}}e^{-4N_{e}}\int_{N_{e}^{i}}^{N_{e}}dN_{e}^{\prime}\ e^{4N_{e}^{\prime}}\frac{T^{7}}{H}\int d\hat{k}\ \hat{k}^{2}\hat{\eta}(\hat{k},T)$ (4.4)
where $N_{e}$ is chosen in the radiation-dominated era and long after the peak of $I_{\mathrm{GW}}$.
@@
$\rho_{\gamma}=\rho_{\gamma,0}\left(\frac{a_{0}}{a}\right)^{4}\left(\frac{g_{*s,0}}{g_{*s}(N_{e})}\right)^{\frac{4}{3}}\quad,\quad\frac{k(N_{e})}{T(N_{e})}=\frac{k_{0}}{T_{0}}\left(\frac{g_{*s}(N_{e})}{g_{*s,0}}\right)^{\frac{1}{3}}$ (4.6)
where $g_{*s,0}=3.931$ is the present degrees of freedom of the entropy density. $\rho_{\gamma,0}$ is the current photon energy density.
@@
$\begin{split}\Omega_{\mathrm{GW},0}(k_{0})h^{2}=&\frac{2}{\pi^{2}M_{\rm pl}^{2}}\frac{\Omega_{\gamma,0}h^{2}}{\rho_{c}(N_{e})}\frac{g_{*}(N_{e})}{g_{\gamma}}\left(\frac{g_{*s,0}}{g_{*s}(N_{e})}\right)^{\frac{1}{3}}\times\\ &e^{-4N_{e}}\int_{N_{e}^{i}}^{N_{e}}dN_{e}^{\prime}\ e^{4N_{e}^{\prime}}\frac{T^{7}}{H}\left(\alpha(N_{e}^{\prime})\frac{k_{0}}{T_{0}}\right)^{3}\hat{\eta}\left[\alpha(N_{e}^{\prime})\frac{k_{0}}{T_{0}}\left(\frac{g_{*s}}{g_{*s,0}}\right)^{\frac{1}{3}},T\right]\end{split}$ (4.7)
where $\Omega_{\gamma,0}=5.38\times 10^{-5}$ , $g_{\gamma}=2$ and the dimensionless Hubble parameter $h=0.674$ [36]. We set $g_{*s}(N_{e})=g_{*}(N_{e})=106.75$. And we have
$\alpha(N_{e}^{\prime})=\frac{T(N_{e})e^{N_{e}}}{T(N_{e}^{\prime})e^{N_{e}^{\prime}}}$ (4.8)
Note that this function stays constant during the radiation-dominated era since $T(N_{e})\propto e^{-N_{e}}$. So in CI, as the GW are produced during the radiation-dominated era, the GW spectrum can be parametrized in the form of
@@
In WI, most GWs are generated around $N_{\mathrm{peak}}^{\mathrm{GW}}$, occurring after inflation and before the radiation-dominated era, as we have discussed. During this stage, $\alpha(N_{e}^{\prime})$ will slowly evolve over time, so the shape of GW spectrum will differ from $\Omega_{\mathrm{GW},0}^{\rm CI}$, and the peak frequency shifts. However, as we will show in our result, this change is very slight and $f_{\rm peak}^{\rm WI}\simeq f_{\rm peak}^{\rm CI}$, where $f_{\rm peak}^{\rm WI}$ is the peak frequency of freeze-in GW in WI.
@@ 5 Freeze-in production of purely gravitational dark matter
Purely gravitational DM [46, 47, 48] is one of the simplest possibilities for DM, since gravitational interaction is universal and model-independent (at least for sub-Planckian scales). It can be produced by the thermal bath mediated by gravitons, known as the gravitational DM freeze-in and it is a concrete realization of the DM freeze-in mechanism. The production of purely gravitational DM in early universe was carefully studied in CI [49, 50, 51, 52, 53], and was first considered in WI in ref. [54]. In this section, we focus on the correlation between the DM mass and the peak amplitude of the GW spectra as functions of the quartic coupling of inflaton field, applying our results from the last section.
@@
$He^{-3N_{e}}\frac{dn_{\chi}e^{3N_{e}}}{dN_{e}}=\beta\frac{T^{8}}{M_{\rm pl}^{4}}$ (5.6)
@@
$Y_{\chi}(N_{e})=\frac{45}{2\pi^{2}g_{*s}}\frac{e^{-3N_{e}}}{T^{3}(N_{e})}\int_{N_{e}^{i}}^{N_{e}}I_{\chi}(N_{e}^{\prime})dN_{e}^{\prime}$ (5.9)
We show the evolution of $Y_{\chi}$ in figure 3. The evolution of the Hubble parameter $H$, the plasma temperature $T$ and the parameter $Q$ are also given in figure 3. During the early stage of WI, although the inflaton energy dominates the universe, the thermal bath is still sustained with a nearly constant temperature and DM gradually accumulates. After the inflation stage, the universe transits to a radiation-dominated phase, where $T\propto e^{-N_{e}}$ and $H\propto T^{2}$.
@@
$\Omega_{\mathrm{DM},0}h^{2}=\frac{m_{\chi}Y_{\chi}(N_{e})s_{0}}{\rho_{c,0}}h^{2}\equiv 0.12$ (5.10)
@@
In figure 4, to satisfy the observed DM relic density, we show the correlation between DM mass and the peak amplitude of GW spectra for different $Q_{\rm ini}$ and dissipation terms. For each point on the curves, we have checked that the condition $m_{\chi}\ll T$ is always satisfied at $N_{e}=N_{\rm peak}^{\rm DM}$, when most DM is produced. Compared with the case $\Upsilon\propto T^{3}$, it can be seen that varying $Q_{\rm ini}$ has a stronger impact on both the predicted DM mass and the peak amplitude of the GW spectra in the case of $\Upsilon\propto T$. Moreover, the DM mass is significantly more sensitive to the coupling $\lambda$ than the GW peak amplitude is. And as the DM mass increases, the peak amplitude of the GW spectra decreases in all cases studied.
@@
Table 1: Model predictions for WI with $\Upsilon\propto T$: scalar spectral index $n_{s}$, tensor-to-scalar ratio $r$, peak frequency $f_{\mathrm{peak}}$, peak amplitude $\Omega_{\mathrm{peak}}h^{2}$, and DM mass $m_{\chi}$
@@
Table 2: Model predictions for WI with $\Upsilon\propto T^{3}$: scalar spectral index $n_{s}$, tensor-to-scalar ratio $r$, peak frequency $f_{\mathrm{peak}}$, peak amplitude $\Omega_{\mathrm{peak}}h^{2}$, and DM mass $m_{\chi}$
@@ 6 Conclusions
Motivated by the recent WI model with solely Standard Model gauge interactions [1], we have investigated the GW spectrum in the high frequency regime generated via the freeze-in mechanism in two types of WI models. We find that in WI, the freeze-in production of GW shows similar features to the DM freeze-in production, with both being enhanced compared with CI. Moreover, the shapes of GW spectra and the peak frequencies ($\sim 10^{10}\penalty 10000\ \rm Hz$) are nearly identical across different WI models, while the peak amplitude of the GW spectra today can be of the order $10^{-13}-10^{-12}$. We also studied the graviton-portal DM freeze-in scenario within two classes of WI framework. Under the constraint of reproducing the observed DM relic abundance, we have identified a correlation between the DM mass and the quartic coupling of the inflation field, and determined the corresponding peak amplitude of the GW signal. Notably, the predicted DM mass exhibits significant variation across different models. These findings provide a theoretical basis for probing such WI scenarios and DM production mechanisms through future GW observations.
@@ Acknowledgments
This work is supported by the National Natural Science Foundation of China (NNSFC) under Grant No.12475111. and No.12205387.
@@
While this work motivated by recent WI with solely Standard Model gauge interactions [1] was nearing completion, ref. [27] appeared on arXiv, which studied the thermal gravitons from warm inflation.

### v4

@@ 4 Gravitational waves spectrum
withk^=k/T\ $\hat${k}=k/T. The $\hat{\eta}$ function takes into account all the Standard Model contributions and has been calculated in refs. [36, 37] using the leading-log approximation. The calculation including hard thermal loop resummation has been done in ref. [38]. Here we only consider the results of leading-log approximation [36],
@@
$\rho_{\mathrm{GW}}(N)=\frac{2}{\pi^{2}M_{\rm pl}^{2}}e^{-4N}\int_{N_{0}}^{N}dN^{\prime}\ e^{4N^{\prime}}\frac{T^{7}}{H}\int d\hat{k}\ \hat{k}^{2}\hat{\eta}(\hat{k},T)$ (4.4)
where $N$ is chosen in the radiation-dominated era and long after the peak of $I_{\mathrm{GW}}$.
@@
$\rho_{\gamma}=\rho_{\gamma,0}\left(\frac{a_{0}}{a}\right)^{4}\left(\frac{g_{*s,0}}{g_{*s}(N)}\right)^{\frac{4}{3}}\quad,\quad\frac{k(N)}{T(N)}=\frac{k_{0}}{T_{0}}\left(\frac{g_{*s}(N)}{g_{*s,0}}\right)^{\frac{1}{3}}$ (4.6)
where $g_{*s,0}=3.931$ [52] is the present degrees of freedom of the entropy density. $\rho_{\gamma,0}$ is the current photon energy density.
@@
$\begin{split}\Omega_{\mathrm{GW},0}(k_{0})h^{2}=&\frac{2}{\pi^{2}M_{\rm pl}^{2}}\frac{\Omega_{\gamma,0}h^{2}}{\rho_{c}(N)}\frac{g_{*}(N)}{g_{\gamma}}\left(\frac{g_{*s,0}}{g_{*s}(N)}\right)^{\frac{1}{3}}\times\\ &e^{-4N}\int_{N_{0}}^{N}dN^{\prime}\ e^{4N^{\prime}}\frac{T^{7}}{H}\left(\alpha(N^{\prime})\frac{k_{0}}{T_{0}}\right)^{3}\hat{\eta}\left[\alpha(N^{\prime})\frac{k_{0}}{T_{0}}\left(\frac{g_{*s}}{g_{*s,0}}\right)^{\frac{1}{3}},T\right]\end{split}$ (4.7)
where $\Omega_{\gamma,0}=5.38\times 10^{-5}$ , $g_{\gamma}=2$ and the dimensionless Hubble parameter $h=0.674$ [52]. We set $g_{*s}(N)=g_{*}(N)=106.75$. And we have defined
$\alpha(N^{\prime})\equiv\frac{T(N)e^{N}}{T(N^{\prime})e^{N^{\prime}}}$ (4.8)
Note that this function stays constant during the radiation-dominated era since $T(N^{\prime})\propto e^{-N^{\prime}}$. So in CI, as the GW are produced during the radiation-dominated era, the GW spectrum can be parametrized in the form of
@@
In WI, most GWs are generated around $N_{\mathrm{peak}}^{\mathrm{GW}}$, occurring after inflation and before the radiation-dominated era, as we have discussed. During this stage, $\alpha(N^{\prime})$ will slowly evolve over time, so the shape of GW spectrum will differ from $\Omega_{\mathrm{GW},0}^{\rm CI}$, and the peak frequency shifts. However, as we will show in our result, this change is very slight and $f_{\rm peak}^{\rm WI}\simeq f_{\rm peak}^{\rm CI}$, where $f_{\rm peak}^{\rm WI}$ is the peak frequency of freeze-in GW in WI.
@@ 5 Freeze-in production of purely gravitational dark matter
Purely gravitational DM [62, 63, 64] is one of the simplest possibilities for DM, since gravitational interaction is universal and model-independent (at least for sub-Planckian scales). It can be produced by the thermal bath mediated by gravitons, known as the gravitational DM freeze-in and it is a concrete realization of the DM UV freeze-in mechanism. The production of purely gravitational DM in early Universe was carefully studied in CI [65, 66, 67, 68, 69], and was first considered in WI in ref. [70]. In this section, we focus on the correlation between the DM mass and the peak amplitude of the GW spectra as functions of the quartic coupling of inflaton field, applying our results from the last section.
@@
$He^{-3N}\frac{dn_{\chi}e^{3N}}{dN}=\beta\frac{T^{8}}{M_{\rm pl}^{4}}$ (5.6)
@@
$Y_{\chi}(N)=\frac{45}{2\pi^{2}g_{*s}}\frac{e^{-3N}}{T^{3}(N)}\int_{N_{0}}^{N}I_{\chi}(N^{\prime})dN^{\prime}$ (5.9)
We show the evolution of $Y_{\chi}$ in figure 3. The evolution of the Hubble parameter $H$, the plasma temperature $T$ and the parameter $Q$ are also given in figure 3. During the early stage of WI, although the inflaton energy dominates the Universe, the thermal bath is still sustained with a nearly constant temperature and DM gradually accumulates. After the inflation stage, the Universe transits to a radiation-dominated phase, where $T\propto e^{-N}$ and $H\propto T^{2}$.
@@
$\Omega_{\mathrm{DM},0}h^{2}=\frac{m_{\chi}Y_{\chi}(N)s_{0}}{\rho_{c,0}}h^{2}\equiv 0.12$ (5.10)
@@
Figure 5: The correlation between DM mass and the peak amplitude of GW spectra for different $Q_{\rm ini}$ and dissipation terms, with $\Omega_{\rm peak}h^{2}$ on the y-axis, $m_{\chi}$ on the x-axis, and $\lambda$ color-coded.
In figure 4 and figure 5, to satisfy the observed DM relic density, we show the correlation between DM mass and the peak amplitude of GW spectra for different $Q_{\rm ini}$ and dissipation terms. For each point on the curves, we have checked that the condition $m_{\chi}\ll T$ is always satisfied at $N=N_{\rm peak}^{\rm DM}$, when most DM is produced. Compared with the case $\Upsilon\propto T^{3}$, it can be seen that varying $Q_{\rm ini}$ has a stronger impact on both the predicted DM mass and the peak amplitude of the GW spectra in the case of $\Upsilon\propto T$. Moreover, the DM mass is significantly more sensitive to the coupling $\lambda$ than the GW peak amplitude is. And as clearly shown in figure 5, the peak amplitude of the GW spectra decreases with increasing DM mass across all cases studied. Notably, the distinction between the two WI models for $Q_{\rm ini}=1$ becomes more pronounced than that for $Q_{\rm ini}=0.01$.
@@
Table 1: Model predictions for WI with $\Upsilon\propto T$: scalar spectral index $n_{s}$, tensor-to-scalar ratio $r$, peak frequency $f_{\mathrm{peak}}$, peak amplitude $\Omega_{\mathrm{peak}}h^{2}$, and DM mass $m_{\chi}$.
@@
Table 2: Model predictions for WI with $\Upsilon\propto T^{3}$: scalar spectral index $n_{s}$, tensor-to-scalar ratio $r$, peak frequency $f_{\mathrm{peak}}$, peak amplitude $\Omega_{\mathrm{peak}}h^{2}$, and DM mass $m_{\chi}$.
@@ 6 Conclusions
Motivated by the recent development in WI model [1], we investigate two important cosmic relics: GW and DM, in the WI scenarios. We have extended the framework of freeze-in graviton production from CI to WI, computing the resulting GW spectrum today in two well-motivated WI models, characterized by dissipation coefficients $\Upsilon\propto T$ and $\Upsilon\propto T^{3}$. We find that the freeze-in production of GW is enhanced in WI, similar to the case of UV freeze-in DM studied in previous literature. Moreover, the shapes of GW spectra and the peak frequencies ($\sim 10^{10}\penalty 10000\ \rm Hz$) are nearly identical across different WI models, while the peak amplitude of the GW spectra today can be of the order $10^{-13}-10^{-12}$. We also studied graviton-portal DM within WI framework, as a specific and simple possibility of UV freeze-in DM. Under the constraint of reproducing the observed DM relic abundance, we have identified a correlation between the DM mass and the quartic coupling of the inflation field, and determined the corresponding peak amplitude of the GW signal. Notably, the predicted DM mass exhibits significant variation across different models and parameters.
In summary, our work explores the connections between GWs, DM, and the microscopic dynamics of WI. We observe that different WI models, characterized by different dissipation coefficients and hence different background evolution, lead to distinguishable GW signals and DM relic abundances. This suggests that future observations combining high-frequency GW detection and DM phenomenology could offer a potential approach to probing the detailed physics of WI models.
@@ Acknowledgments
This work is supported by the National Natural Science Foundation of China (NNSFC) under Grant No.12475111. and No.12205387, and the Fundamental Research Funds for the Central Universities, Sun Yat-sen University.
@@
While this work, motivated by the recent development in WI [1], was nearing completion, ref. [42] appeared on arXiv, which studied the thermal gravitons from warm inflation.
Appendix A UV freeze-in dark matter in warm inflation
In this appendix, we briefly review the general aspects of UV freeze-in DM in WI.
In the framework of UV freeze-in DM [43], the hidden and visible sectors are coupled only via non-renormalisable operators.The evolution of the number density of DM is governed by the following Boltzmann equation.
$\frac{dn_{\rm DM}}{dt}+3Hn_{\rm DM}=\mathcal{C}\frac{T^{2n+4}}{\Lambda^{2n}}$ (A.1)
The source term on the right-hand side represents the effective DM-radiation interaction arising from a non-renormalisable operator of mass dimension $n+4$, with $\Lambda$ being the cutoff scale of the effective description and $\mathcal{C}$ a dimensionless coefficient. For the example of purely gravitational DM discussed in section. 5, $\Lambda$ is just the Planck mass and $n=2$, $\mathcal{C}=3096\zeta_{3}^{2}/5\pi^{3}$.
Using $dN\equiv Hdt$, the solution of eq. (A.1) reads:
$n_{\rm DM}(N)=e^{-3N}\int_{N_{0}}^{N}I_{\rm DM}(N^{\prime})dN^{\prime}$ (A.2)
where
$I_{\rm DM}(N)\equiv\mathcal{C}e^{3N}\frac{T^{2n+4}(N)}{\Lambda^{2n}H(N)}$ (A.3)
is the production rate of the comoving DM number density: $I_{\rm DM}=d(e^{3N}n_{\rm DM})/dN$.
According to eq. (A.2), the DM number density depends on the evolution of temperature and Hubble parameter, and therefore on the specific WI model. However, an overarching conclusion can still be derived as follows.
During the inflationary phase, both the inflaton and radiation undergo slow-roll evolution, causing the temperature $T(N)$ and the Hubble parameter $H(N)$ to remain approximately constant. Consequently, $I_{\rm DM}(N)$ becomes an exponentially increasing function. However, after inflation ends and the radiation-dominated era begins, it becomes exponentially decreasing, as $T(N)\propto e^{-N}$ and $H(N)\propto e^{-2N}$. Therefore, the function $I_{\rm DM}(N)$ is expected to exhibit a sharp peak at a specific e-fold, $N_{\rm peak}^{\rm DM}$, determined by $dI_{\rm DM}(N)/dN=0$. Explicitly, this condition reads:
$\left.\left[3+(2n+4)\frac{d\ln{T(N)}}{dN}-\frac{d\ln H(N)}{dN}\right]\right|_{N=N_{\mathrm{peak}}^{\rm DM}}=0$ (A.4)
This property allows us to estimate the DM number density at some late times after $N_{\rm peak}^{\rm DM}$. We obtain:
$n_{\rm DM}(N)\simeq e^{-3N}I_{\rm DM}(N_{\rm peak}^{\rm DM})\Delta N_{\mathrm{peak}}^{\rm DM}\quad(N>N_{\mathrm{peak}}^{\rm DM})$ (A.5)
with $\Delta N_{\mathrm{peak}}^{\rm DM}\gtrsim 1$ being the half-width of $I_{\rm DM}(N)$.
The DM yield, $Y_{\rm DM}\equiv n_{\rm DM}/s$, can then be estimated [44],
$Y_{\rm DM}(N)\simeq\frac{45}{2\pi^{2}g_{*s}}\frac{e^{3(N_{\mathrm{peak}}^{\rm DM}-N)}}{\Lambda^{2n}T^{3}(N)}\Delta N_{\mathrm{peak}}^{\rm DM}\times\frac{T^{2n+4}(N_{\mathrm{peak}}^{\rm DM})}{H(N_{\mathrm{peak}}^{\rm DM})}\quad(N>N_{\mathrm{peak}}^{\rm DM})$ (A.6)
To better understand this result, we compare it with the final DM yield produced during radiation dominated epoch with the reheating temperature $T_{RH}\equiv T(\epsilon_{H}=2)$:
$Y_{\rm DM}^{\mathrm{RH}}\simeq\frac{1}{4\sqrt{\pi}}\left(\frac{45}{\pi^{2}g_{*}}\right)^{\frac{3}{2}}\frac{1}{2n-1}\frac{M_{\rm pl}T_{\mathrm{RH}}^{2n-1}}{\Lambda^{2n}}$ (A.7)
In WI, DM production begins at the onset of inflation and always peaks before radiation domination. It follows that the DM yield is expected to be significantly larger than that which would have accumulated from the radiation-dominated epoch alone. This enhancement is then quantified by the ratio:
$\frac{Y_{\rm DM}(N)}{Y_{\rm DM}^{\mathrm{RH}}}\simeq(2n-1)\frac{I_{\rm DM}(N_{\mathrm{peak}}^{\rm DM})}{I_{\rm DM}(N_{\mathrm{RH}})}\Delta N_{\mathrm{peak}}^{\rm DM}$ (A.8)
Figure 6: The evolution of $I_{\mathrm{DM}}$ and $I_{\mathrm{GW}}$ in WI with dissipation term $\Upsilon\propto T^{3}$ and $\Upsilon\propto T$, respectively. All curves are normalized by their peak values. The blue line corresponds to $Q_{\mathrm{ini}}=0.01$ and the red line corresponds to $Q_{\mathrm{ini}}=1$. Vertical lines mark key transitions: the black dashed line indicates inflation termination ($\epsilon_{H}=1$), while the colored dashed lines show the start of radiation domination ($\epsilon_{H}=2$) for each $Q_{\rm ini}$ values. Note that $I_{\mathrm{DM}}$ and $I_{\mathrm{GW}}$ both sharply peaked at some e-folding numbers in the post-inflationary epoch, prior to the onset of radiation domination.
This comparison is equivalent to comparing the DM yield in WI with that in CI under the same reheating temperature $T_{\rm RH}$, since in the CI scenario, DM production occurs predominantly during the radiation-dominated era. And given $I_{\rm DM}(N_{\mathrm{peak}}^{\rm DM})\gg I_{\rm DM}(N_{\rm RH})$ with $\Delta N_{\mathrm{peak}}^{\rm DM}\gtrsim 1$,it follows that the relic abundance is significantly enhanced in WI compared to CI.
Figure 6 shows the evolution of $I_{\mathrm{DM}}$ and $I_{\mathrm{GW}}$ in WI for both $\Upsilon\propto T^{3}$ and $\Upsilon\propto T$, under both $Q_{\mathrm{ini}}=0.01$ and $Q_{\mathrm{ini}}=1$. We set $n=2$ since this is the case of purely gravitational DM as we have discussed in section. 5. All curves are normalized to their respective peak values. The blue and red solid lines represent the cases with $Q_{\mathrm{ini}}=0.01$ and $Q_{\mathrm{ini}}=1$, respectively. The black dashed line marks the end of inflation, while the blue and red dashed lines indicate the onset of radiation domination for each scenario.
As clearly shown in this figure and section. 3, graviton production bears notable similarities to the UV freeze-in of DM. Most graviton energy and DM particles are generated prior to the radiation-dominated epoch. The ratio of their respective production enhancements can be estimated on the order of $I_{X}(N^{X}_{\rm peak})/I_{X}(N_{RH})$, where $X$ stands for DM or GW.